OpenStudy (anonymous):

Solve 2 - a-4/a+3 = a^2-2/a+3 a = ?

OpenStudy (anonymous):

\[2-\frac{a-4}{a+3}=\frac{a^2-2}{a+3}\]?

OpenStudy (anonymous):

satellite73 please help me :) im new at this and i heard you were very smart :)

OpenStudy (anonymous):

Question: As a salesperson at Snow Sports, Micheal receives a monthly base pay plus commission on all that he sells. If he sells $600 worth of merchandise in one month , he is paid $344. If he sells $1000 worth of merchandise, he is paid $440. Find Michael's salary when he sells $2,200 worth of merchandise.

OpenStudy (anonymous):

@jackie your job is to find the equation for the line between the points (600,344) and (1000,440) once you get the equation plug in 2200 and see what you get.

OpenStudy (anonymous):

@jazi1 you job is to add the expressions on the left hand side first. then since the denominators will be the same set the numerators = and solve.

OpenStudy (anonymous):

jackie did you post your question?

OpenStudy (anonymous):

we can work it out if you like

OpenStudy (anonymous):

yes i did :) and please & thnk you love <3

myininaya (myininaya):

1128

OpenStudy (anonymous):

satellite 73 would you work it out please

myininaya (myininaya):

hey jackie i can help you out at your thread i think satellite was trying to help someone here

OpenStudy (anonymous):

yes please :) im sorry satellite & jazil im sorry please calm down .